OpenStudy (anonymous):

Let f(x)=(x+12) * ln(x+1) for x>-1. Then f(x)is concave upward on the interval (a,infinity) where a=___

OpenStudy (anonymous):

Harold and Kumar?

OpenStudy (anonymous):

Just kumar

OpenStudy (anonymous):

Harold is in white castle right now

OpenStudy (anonymous):

Can Kumar be of assistance?

OpenStudy (anonymous):

Depends on the situation.

OpenStudy (anonymous):

The situation above ^^^ lols

OpenStudy (across):

\[f(x)=\ln(x+1)(x+12)\]\[f'(x)=\frac{x+12}{x+1}+\ln(x+1)\]\[f''(x)=\frac{x-10}{(x+1)^2}.\]It follows that \(f''(x)>0\) when \(x>10\).

OpenStudy (anonymous):

what rules did you use?

OpenStudy (across):

I just differentiated the function twice.

OpenStudy (anonymous):

Yeah, I know. Could you be more specific?

OpenStudy (across):

\[\frac{d}{dx}\left[\ln(x+1)(x+12)\right]=\frac{d}{dx}\left[\ln(x+1)\right](x+12)+\ln(x+1)\frac{d}{dx}\left[(x+12)\right]\]\[=\frac{x+12}{x+1}\frac{d}{dx}\left[x+1\right]+\ln(x+1)\cdot1=\frac{x+12}{x+1}\cdot1+\ln(x+1)\]\[\frac{x+12}{x+1}+\ln(x+1)\]Then do that a second time.

OpenStudy (across):

That's the product rule.

OpenStudy (across):

With the chain rule applied to the parentheses.

OpenStudy (anonymous):

ok. thanks. you're really pretty by the way. :)